please show all steps

8.2.2 Use Table V in Appendix A to find the values of the following percentiles: (Express your answers to 3 decimal places.) Part 1 to.025,15 the absolute tolerance is+/-0.001
Answer #1

Solution :

Degrees of freedom = df = 15

\alpha = 0.05

\alpha / 2 = 0.05 / 2 = 0.025

Using t table,

t\alpha /2,df = t0.025,15 = 2.131

t0.025,15 = 2.131